OpenStudy (anonymous):

Simplify 7 log3 k + 6 log3 m – 9 log3 n

OpenStudy (mertsj):

\[a \log_{3}b= \log_{3}b ^{a} \]

OpenStudy (mertsj):

Rewrite each term using that property

OpenStudy (mertsj):

What did you get?

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

do you mean to write as one log?

OpenStudy (kropot72):

Hint \[7\log_{3} k=\log_{3} k ^{7}\]

OpenStudy (anonymous):

another hint: log(MN) = logM + logN

OpenStudy (anonymous):

one more hint: log(M/N) = logM - logN

OpenStudy (anonymous):

I don't understand what I need to do.

OpenStudy (anonymous):

here's an example: log(5) + log(3) = log(5*3) = log(15) i started with the sum of two logs and eventually i ended up with 1 log. that is what you need to do with the expression they gave you.